# Satisfaction-Aware Incentive Scheme for Federated Learning in Industrial Metaverse: DRL-Based Stackbelberg Game Approach

### v1

@@ Abstract
Industrial Metaverse leverages the Industrial Internet of Things (IIoT) to integrate data from diverse devices, employing federated learning and meta-computing to train models in a distributed manner while ensuring data privacy. Achieving an immersive experience for industrial Metaverse necessitates maintaining a balance between model quality and training latency. Consequently, a primary challenge in federated learning tasks is optimizing overall system performance by balancing model quality and training latency. This paper designs a satisfaction function that accounts for data size, Age of Information (AoI), and training latency. Additionally, the satisfaction function is incorporated into the utility functions to incentivize node participation in model training. We model the utility functions of servers and nodes as a two-stage Stackelberg game and employ a deep reinforcement learning approach to learn the Stackelberg equilibrium. This approach ensures balanced rewards and enhances the applicability of the incentive scheme for industrial Metaverse. Simulation results demonstrate that, under the same budget constraints, the proposed incentive scheme improves at least 23.7% utility compared to existing schemes without compromising model accuracy.
@@
Metaverse, Federated Learning, Age of Information, Incentive Scheme, Stackelberg Game.
@@ I Introduction
Metaverse is undoubtedly one of the most popular and promising intelligent applications [1, 2, 3], creating a collective virtual shared space that merges digital and physical realities, enabling users to interact with each other and digital environments in immersive, and real-time settings. Metaverse can also be combined with the Industrial Internet of Things (IIoT), giving rise to industrial Metaverse. For example, NVIDIA [4] developed an open platform called Omniverse, which supports multi-user real-time 3D simulation and visualization of physical properties in a shared virtual space for industrial applications such as automotive design. The meta-computing plays a crucial role in integrating distributed computing resources within the IIoT to build industrial Metaverse. It enables distributed data processing, storage, and computation across heterogeneous systems, bridging the gap between virtual and physical spaces [5]. This provides a seamless platform for industrial Metaverse, delivering a realistic, persistent, and smooth interaction experience for users.
However, traditional distributed learning approaches may face certain limitations due to the large number of devices and data in industrial Metaverse enabled by meta-computing. First, IIoT nodes need to share data for model training, which increases the risk of privacy leakage. Second, the large volume of data transmission may result in latency issues. Fortunately, Federated Learning (FL) [6, 7] can effectively address these problems. FL is a distributed learning scheme proposed by Google to optimize global machine learning models without moving data out of local devices. In the industrial Metaverse, which uses meta-computing and incorporates FL, each IIoT node participates in training a shared AI model using its own dataset, and then uploads its local model to the server to build a new global model [8, 9], thus achieving the goal of building a high-quality Metaverse while ensuring privacy.
To ensure an immersive experience in industrial Metaverse, high-quality global models must be developed. This requires vast amounts of real-time sensing data and significant computational resources from nodes [10]. Furthermore, achieving an immersive experience in the Metaverse demands not only high-quality models but also low-latency interactions. Balancing the reduction of latency with the enhancement of quality is crucial to maintaining the overall performance of Metaverse applications. However, due to the large number of nodes and the complex structure of the industrial Metaverse, designing a methodology that effectively balances quality and latency presents a significant challenge.
To address these challenges, we define a satisfaction function based on data size, Age of Information (AoI) [11], and latency, aiming to balance training latency and model quality. This satisfaction function is then incorporated into the utility function, transforming the utility optimization problem into a Stackelberg game. However, traditional methods such as backward induction, convex optimization, and nonlinear programming require lots of participant information to solve for game equilibrium, which is not practical for industrial Metaverse. The nodes in the industrial Metaverse are not only large-scale but also operate independently. To preserve privacy, these participants are often unwilling to share their private information, and thus lack the a priori information required by traditional methods. Fortunately, Deep Reinforcement Learning (DRL) can learn optimal strategies based on experience alone, eliminating the need for prior information. We propose a DRL-based method to obtain game equilibrium effectively, ensuring participant privacy. The main contributions of this paper are as follows:
  • We design a new meta-computing framework based on FL to provide dynamic resource scheduling and optimization solutions for the industrial Metaverse. The framework incorporates an FL incentive scheme designed for the industrial Metaverse, enabling resource allocation and optimization while ensuring the requirements of the demand side, thereby improving learning efficiency.
@@
  • We introduce a new metric named “Satisfaction” that balances training latency and model quality in the industrial Metaverse. Focusing on the real-time relevance and quality of data, we use factors such as data size, AoI, and service latency as the key factors to measure the contribution of nodes in the industrial Metaverse. We design a satisfaction-aware incentive scheme that constructs utility functions for nodes and servers to incentivize resource sharing among nodes.
@@
  • We transform the optimization problem involving two utility functions into a Stackelberg game model and prove the uniqueness of the Stackelberg Equilibrium (SE). We then use the DRL algorithm to learn the SE without requiring private information from other agents, relying solely on past experiences.
@@ II Related Works
II-A Quality of Metaverse Applications
The definition of quality in the Metaverse varies across different services. The authors in [12] defined the quality of perception in VR as a measure of users’ subjective feelings about their immersion in the virtual world and designed a double Dutch auction mechanism to determine optimal pricing and allocation rules, facilitating fast trades of VR services between users and providers. The authors in [13] proposed a novel metric called Meta-Immersion, which combines objective KPIs with subjective perceptions of Metaverse users, allowing for distinct expressions in different virtual scenarios. They also developed an attention-aware rendering capacity allocation scheme to enhance QoE. The authors in [14] used collected synchronization data and the value decay rate of the digital twin to determine synchronization strength, maximizing its payoff, and employed a dynamic Stackelberg game to obtain optimal control. The authors in [15] proposed a novel metric called Age of Migration Task (AoMT) to quantify the task freshness of Vehicle Twins’ migration, aiming to ensure the quality of migration within the Metaverse, and designed an AoMT-based contractual model to effectively incentivize Vehicle Twins’ migration. Additionally, most existing Metaverse QoE evaluation studies primarily focus on subjective aspects such as video [16], audio [17], 3D models, and AR/VR QoE [18].
However, Metaverse immersion is primarily influenced by image quality and response speed [19, 20], while model quality and latency in FL directly affect virtual image quality and response speed in the Metaverse [21], which leads to reduced immersion. Therefore, the quality metrics defined in the above studies are insufficient for scenarios requiring high immersion, particularly those involving the integration of the industrial Metaverse with FL. Immersive experiences in the industrial Metaverse require a high degree of real-time interaction, which exacerbates the efficiency issues in FL. Thus, identifying quality evaluation methods applicable to the integration of the industrial Metaverse and FL is the focus of this paper.
II-B Incentive Schemes for FL
Node contribution evaluation: Node contribution evaluation in FL focuses on the extent to which each participating node contributes to the overall learning process and outcome, enabling FL to achieve higher performance with minimal rewards [22]. The authors in [23] used historical learning records to estimate nodes’ learning quality and an exponential forgetting function to assign weights, designed quality-aware incentives and model aggregation methods to improve learning effectiveness. The authors in [24] designed the quality-aware node selection framework AUCTION, which included factors such as data size, data quality, and learning budget within nodes that influence learning quality. These factors were encoded into an attention-based neural network to enhance strategy performance. The authors in [25] introduced reputation as a node evaluation metric to improve the reliability of FL tasks in mobile networks, resulting in a reliable worker selection scheme. Shapley values are also used in FL to measure how much a participant contributes to the overall model training quality. The authors in [26] leveraged Shapley value calculation to build FedCoin, a blockchain-based peer-to-peer payment system that ensures realistic and fair profit distribution.
Incentive Schemes: Auction-based schemes are commonly applied in FL due to their simplicity of construction. The authors in [27] proposed an incentive mechanism named FMore, designed for multidimensional procurement auctions, to encourage more low-cost, high-quality edge nodes to participate in learning. The authors in [23] proposed an FL system called FAIR, where reverse auctions were modeled to encourage quality users to participate in FL. The Stackelberg game, a sequential model in game theory, is suitable for incentive design in FL. The authors in [28] proposed an FL-based crowdsourcing framework using incentive-driven interactions and analyzed it with a two-stage Stackelberg game. The authors in [29] designed a dynamic incentive scheme based on the Stackelberg game to adaptively adjust node selection. Reinforcement Learning (RL) adapts to the environment and optimizes the strategy through continuous iteration in a dynamic setting [30]. The authors in [31] proposed a DRL-based mechanism to optimize pricing for servers and training policies for edge nodes, incentivizing participation. Additionally, several studies have integrated game theory with DRL. The authors in [32] formulated the interaction between edge servers and data centers as a multi-leader multi-follower Stackelberg game, employing the MADDPG [33] algorithm to achieve optimal strategies. The authors in [34] modeled providers in the Metaverse as a Stackelberg game and used MALPPO to optimize Vehicular Twin migration. Several papers also proposed incentives for FL in the Metaverse. The authors in [35] proposed a decentralized FL framework with privacy preservation and an age-based model to incentivize IIoT Metaverse data sensing. The authors in [36] proposed a Metaverse optimization framework minimizing energy, time and accuracy trade-offs, with a resource allocation algorithm for device contributions.
@@ II-B Incentive Schemes for FL
Figure 1: A meta-computing framework based on FL for industrial Metaverse.The framework enables resource scheduling and task management for IIoT devices, where the server employs a Satisfaction-aware incentive mechanism to coordinate nodes for efficient task execution.
However, none of these works identify the potential factors affecting FL quality related to the freshness of information, nor do they consider the overall budgetary investment. Therefore, designing an FL incentive mechanism based on model quality that meets the needs of the industrial Metaverse is another focus of this paper.
@@ III System Model
In order to promote efficient task processing and decision making in the industrial Metaverse, we design a meta-computing framework based on FL in Fig. 1. The device management module contains multiple edge nodes. Its primary purpose is to collect data from production devices, integrate the computational, storage, and communication resources of the edge nodes, map these resources to the server, convert them into objects that can be easily accessed by the resource scheduler, and then train FL models locally based on the incentive scheme developed by the task manager. The resource scheduler module contains several virtual edge nodes that constantly monitor changes in the configuration details of physical nodes, simulate possible states of the nodes, and dynamically perform resource optimization. The task management module, located on the server, accepts requests from users, decomposes tasks, and designs incentive schemes based on their conditional constraints. The zero-trust computing management module performs global aggregation for FL via the blockchain, and the identity and access management module ensures users have the appropriate permissions to access the data. Since the zero-trust computing management and identity and access management modules are not the focus of this paper, detailed discussions are omitted. Comprehensive analyses of these modules are available in [5]. The main notations used in this paper are shown in Table I.
@@
$\begin{split}\omega*=\arg\min F(\omega).\end{split}$ (3)
Due to the inherent complexity of many machine learning models, it is hard to find a closed-form solution. Therefore, it is often solved using gradient descent techniques [37]. Given that IIoT nodes may be reluctant to provide new sensing data for time-sensitive FL tasks in the industrial Metaverse, reliable incentives are greatly needed to encourage users to share new sensing data, which is discussed in the next section.
IV Satisfaction: Quality Control for Industrial Metaverse
The structure of the industrial Metaverse is complex, containing multiple types of industrial devices and involving a large number of industrial nodes, where the trade-off between low latency and high model quality must be carefully considered. Among them, latency directly affects screen lag and response time in virtual environments. High latency causes the screen to fail to update in time, resulting in users experiencing lags or delays when interacting with the environment [19, 20]. In other words, user interaction in the Metaverse is highly dependent on low latency, and any delay will significantly reduce user satisfaction. On the other hand, degraded model quality means that the information or imagery received by users is not realistic enough, which may lead to a lack of synchronization between the virtual environment and the user’s actual behavior. Furthermore, model quality will greatly disturb users’ decisions. With information from a declining quality model, users may make inappropriate responses. Traditional quality metrics in FL do not adequately capture the requirements of the industrial Metaverse. Therefore, we propose a satisfaction metric $G_{i}$, to balance quality and latency. The satisfaction metric $G_{i}$ of the task $i$ is denoted as
@@ IV Satisfaction: Quality Control for Industrial Metaverse
In the context of combining the industrial Metaverse with FL, the final aggregated global model depends on the local models contributed by individual nodes, while the size of the raw data ${D_{i}}$ reflects the node’s contribution to the overall FL training process. ${D_{i}}$ is denoted as ${D_{i}}=\frac{T}{{{\theta_{i}}}}d$, where $d$ is the amount of data collected per unit time period.
@@ V-A Utility Model and Problem Formulation
To meet the quality and latency requirements of industrial Metaverse tasks, we formulate the corresponding utility functions and optimization objectives. Node $i$ is incentivized to handle incoming task requests, and each participating node receives a monetary reward ${R_{i}}$ from the server. Therefore, the utility of node $i$ is the difference between the reward ${R_{i}}$ and the cost ${C_{i}}$ of participating in the FL task. The utility $U_{i}$ can be expressed as
@@
In order to meet the quality and latency requirements of industrial Metaverse tasks, we define the goal of the incentive mechanism as maximizing server utility function $V$ and node utility function $U_{i}$. The optimization problem for the node utility is formulated as follows:

### v2

# Meta-Computing Enhanced Federated Learning in IIoT: Satisfaction-Aware Incentive Scheme via DRL-Based Stackelberg Game

@@ Abstract
The Industrial Internet of Things (IIoT) leverages Federated Learning (FL) for distributed model training while preserving data privacy, and meta-computing enhances FL by optimizing and integrating distributed computing resources, improving efficiency and scalability. Efficient IIoT operations require a trade-off between model quality and training latency. Consequently, a primary challenge of FL in IIoT is to optimize overall system performance by balancing model quality and training latency. This paper designs a satisfaction function that accounts for data size, Age of Information (AoI), and training latency for meta-computing. Additionally, the satisfaction function is incorporated into the utility functions to incentivize nodes in IIoT participation in model training. We model the utility functions of servers and nodes as a two-stage Stackelberg game and employ a deep reinforcement learning approach to learn the Stackelberg equilibrium. This approach ensures balanced rewards and enhances the applicability of the incentive scheme for IIoT. Simulation results demonstrate that, under the same budget constraints, the proposed incentive scheme improves utility by at least 23.7% compared to existing FL schemes without compromising model accuracy.
@@
Industrial Internet of Things, Federated Learning, Age of Information, Incentive Scheme, Stackelberg Game.
@@ I Introduction
The Industrial Internet of Things (IIoT) serves as a core pillar of intelligent manufacturing and digital industrial systems, enabling real-time data collection, analysis, and intelligent decision making by connecting a large number of industrial devices, sensors, and intelligent systems [1, 2, 3]. Meta-computing [4] is an emerging computing paradigm that integrates and optimizes distributed computing resources in IIoT through key modules such as zero-trust computing management, task management, device management, identity and access management, and resource scheduling. These modules enable secure aggregation, dynamic task allocation, and efficient resource orchestration, providing efficient, flexible, and scalable computing services. Federated Learning (FL) [5, 6] is a collaborative AI approach that enables IIoT to optimize global machine learning models without migrating local data from IIoT nodes, thus enhancing the intelligent analytics capabilities of meta-computing while ensuring data privacy [7, 8].
To ensure efficient operation of IIoT, FL in meta-computing must meet key requirements such as efficient resource allocation, high-quality models, and low latency [9]. However, due to the large scale of IIoT devices and the complexity of data structures, FL in meta-computing still faces several challenges.
First, IIoT systems consist of numerous heterogeneous devices with varying computational capabilities and communication bandwidths, while task demands fluctuate dynamically with changes in the production environment. Some low-power devices struggle with high-load computational tasks, whereas high-performance computing nodes may experience bottlenecks due to unbalanced resource allocation, reducing overall system efficiency. Therefore, a key challenge is how to efficiently schedule tasks within limited resources to optimize resource allocation. Second, machine learning models are widely used in IIoT for intelligent prediction, anomaly detection, and automated control. However, high-quality models typically require significant computational resources and longer training times, while IIoT scenarios with large data volumes and complex structures often demand low latency and real-time responsiveness. Moreover, since IIoT nodes operate in distributed and autonomous environments, they are generally unwilling to share private information, further complicating the balance between model quality and latency [9]. Consequently, reducing computational and communication latency while maintaining model quality remains a critical research challenge.
To address these challenges, we propose a novel evaluation metric, Satisfaction, to quantify the contribution of IIoT nodes in the FL process. This metric considers data volume, Age of Information (AoI) [10], and computation latency to assess node value. Specifically, AoI measures model quality, prioritizing newer models in updates, while computation latency evaluates node responsiveness, preventing high-latency devices from degrading overall training efficiency. Building on this metric, we develop a Satisfaction-Aware Incentive Mechanism that constructs utility functions for both nodes and the server, dynamically adjusts budget allocations, and incentivizes resource sharing. This mechanism encourages high-data-quality and high-efficiency nodes to actively participate in training. Furthermore, we model the incentive mechanism as a Stackelberg Game, where IIoT nodes act as followers responding to the server’s leader decisions. Unlike traditional game-theoretic approaches, we employ Deep Reinforcement Learning (DRL) to learn optimal strategies from historical experience, eliminating the need for private information from participating agents. Our main contributions are summarized as follows.
  • We propose Meta-Computing Enhanced Federated Learning (MEFL) for dynamic computation scheduling in IIoT. MEFL leverages meta-computing resource scheduling to optimize FL computational resource allocation, employs task management to dynamically orchestrate FL training tasks, utilizes device management to coordinate distributed computing nodes, integrates zero-trust computing management to improve data privacy protection, and incorporates identity and access management for authentication and access control, ensuring secure resource access. MEFL empowers IIoT with efficient, secure, and adaptive computing capabilities, enabling intelligent decision-making and optimized resource utilization in distributed environments.
@@
  • We introduce a new metric named “Satisfaction” to balance training latency and model quality. By considering data size, AoI, and computation latency, we measure the contribution of IIoT nodes and develop a Satisfaction-Aware incentive scheme, constructing utility functions for both nodes and the server to encourage resource sharing.
@@
  • We transform the optimization problem involving two utility functions into a Stackelberg Game Model and prove the uniqueness of the Stackelberg Equilibrium (SE). We then employ a DRL-based algorithm to learn the SE without requiring private information from other agents, enabling optimal strategy learning purely from historical experience.
@@ II Related Works
II-A FL in IIoT
FL has been widely applied in IIoT to enable decentralized intelligence while preserving data privacy. However, the heterogeneous nature of IIoT devices poses significant challenges for efficient resource scheduling. The authors in [11] proposed a semi-decentralized federated edge learning (SD-FEEL) framework that optimizes device association, resource allocation, and edge server placement to minimize training loss and improve accuracy within a limited cost budget. The authors in [12] proposed a client scheduling scheme for a semi-decentralized FL (SD-FL) framework in IIoT, optimizing client-server association to enhance communication reliability and training efficiency while minimizing global training loss. The authors in [13] proposed a FL-based Cyber Threat Intelligence framework (FL-CTIF) that enhances IIoT security by leveraging information fusion to improve cyberattack detection accuracy while reducing training rounds and CPU consumption. The authors in [14] proposed a decentralized federated learning framework for edge-enabled smart manufacturing, optimizing industrial data processing with inexact ADMM algorithms to enhance low-latency and secure decision-making while improving response time and maintaining accuracy. The authors in [15] proposed a DT-empowered IIoT architecture optimized with FL, integrating deep reinforcement learning for device selection and an asynchronous FL scheme to address device heterogeneity. The authors in [16] proposed a framework FDEI integrating federated learning with DT-enabled IIoT to enhance service quality and trustworthiness. The authors in [17] proposed an energy-efficient federated learning framework for digital twin-enabled IIoT, where IIoT devices dynamically select training methods based on environmental conditions. The authors in [18] introduced a meta-computing-driven vertical FL algorithm to address data incompleteness in heterogeneous IIoT devices, enabling collaborative intelligence without requiring shared feature spaces.
However, existing studies mainly focus on integrating DT with FL while lacking effective mechanisms for dynamic resource scheduling.
II-B Incentive Schemes in IIoT
FL relies on distributed nodes to jointly participate in model training, and the performance of nodes directly affects the training latency and model quality [19]. Without an effective incentive scheme, selfish nodes might not provide sufficient resources or may be reluctant to join the federated learning task [20]. Therefore, designing an incentive scheme that fairly evaluates node contributions and encourages high-quality participation is crucial. The authors in [21] proposed the Model Value Transfer Incentive (MVTI) to enhance federated learning incentives for AIoT, ensuring secure execution and fair benefit redistribution with smart contracts and IPFS. The authors in [22] proposed BIMFL, a bilevel optimization-based incentive mechanism for federated learning, where the parameter server assigns rewards, and devices decide participation to maximize profit based on rewards and energy costs. The authors in [23] proposed a communication-efficient incentive mechanism for federated learning using Nash bargaining theory, formulating a concurrent bargaining game with a probabilistic greedy client selection algorithm. The authors in [24] proposed a multi-exit-based federated edge learning (ME-FEEL) framework for resource-constrained IIoT, enabling devices with limited computational power to exit early during training, reducing latency and maximizing participation in model aggregation. The authors in [25] proposed a social credit-based incentive mechanism to enhance cooperative D2D transmission across IIoT networks, improving throughput and reducing power consumption.
@@ II-B Incentive Schemes in IIoT
Figure 1: The MEFL framework for IIoT. The framework enables resource scheduling and task management for IIoT devices, where the server employs a Satisfaction-aware incentive mechanism to coordinate nodes for efficient task execution.
However, existing incentive mechanisms primarily focus on reward distribution and participation decisions but lack a comprehensive strategy to balance training latency and model quality. Most approaches fail to incorporate real-time factors such as data freshness and computational efficiency when evaluating node contributions, leading to suboptimal training performance in dynamic IIoT environments.
II-C Game Theoretic and DRL Approaches in IIoT
Federated learning in IIoT involves complex resource allocation challenges. Traditional heuristic-based scheduling methods often fail to adapt to dynamic environments. Game-theoretic models have been explored to address resource competition in hierarchical decision-making scenarios, while DRL has emerged as a promising technique for learning optimal resource allocation policies in complex, dynamic IIoT networks. The authors in [26] proposed a blockchain-based resource trading framework for multi-UAV-assisted IIoT networks, leveraging Stackelberg game modeling and multi-agent DRL to optimize dynamic resource allocation while ensuring security and privacy. The authors in [27] proposed a DRL-based cloud-edge task scheduling model LsiA3CS, leveraging a Markov game and an asynchronous advantage actor-critic (A3C) algorithm to balance computing resources and reduce communication latency in large-scale IIoT networks. The authors in [28] focused on Stackelberg game-driven dynamic pricing and resource allocation, utilizing deep Q-networks (DQN) to approximate optimal strategies for resource pricing and task scheduling in IIoT.
However, existing approaches primarily focus on resource allocation and task offloading but do not explicitly incorporate Stackelberg game-theoretic modeling into FL for hierarchical optimization. Most solutions assume complete system information, which is impractical in privacy-sensitive FL scenarios where devices operate autonomously. Additionally, many DRL-based frameworks lack an incentive-aware mechanism to guide learning policies effectively.
@@ III System Model
In order to promote efficient task processing and decision making in IIoT, we design MEFL, a meta-computing enhanced FL framework, as shown in Fig. 1. The device management module contains multiple edge nodes, whose primary purpose is to collect data from production devices, integrate the computational, storage, and communication resources of the edge nodes, and map these resources to the server. These resources are then converted into objects that can be easily accessed by the resource scheduler, enabling local FL training based on the incentive scheme developed by the task manager. The resource scheduler module contains several digital twins of edge nodes, which continuously monitor changes in the configuration details of physical nodes, simulate their possible states, and dynamically optimize resource allocation. The task management module, located on the server, accepts user requests, decomposes tasks, and designs incentive schemes based on conditional constraints. The zero-trust computing management module performs global aggregation for FL via blockchain, while the identity and access management module ensures users have appropriate permissions to access data. Since the zero-trust computing management and identity and access management modules are not the focus of this paper, detailed discussions are omitted. Comprehensive analyses of these modules are available in [4]. The main notations used in this paper are shown in Table I.
$\begin{split}\omega*=\arg\min F(\omega).\end{split}$ (3)
Due to the inherent complexity of many machine learning models, it is hard to find a closed-form solution. Therefore, it is often solved using gradient descent techniques [29]. Given that IIoT nodes may be reluctant to provide new sensing data for time-sensitive FL tasks, reliable incentives are greatly needed to encourage users to share new sensing data, which is discussed in the next section.
@@
IV Satisfaction: Quality Control for IIoT
The structure of IIoT is complex, containing multiple types of industrial devices and involving a large number of industrial nodes, where the trade-off between low latency and high model quality must be carefully considered in meta-computing enhanced federated learning. Latency plays a critical role in industrial control and decision-making processes. High latency can lead to delayed data synchronization, causing inefficiencies in predictive maintenance, anomaly detection, and real-time monitoring [30]. Meanwhile, model quality determines the accuracy and reliability of decision-making. Degraded model quality means that the data used for decision-making is less accurate or outdated, which may lead to incorrect control actions, reduced production efficiency, or potential safety risks. Furthermore, model quality directly influences the effectiveness of IIoT decision-making processes—low-quality models may lead to suboptimal system adjustments or improper responses to changing industrial conditions. Traditional quality metrics in FL do not adequately capture the requirements of IIoT environments. Therefore, we propose a satisfaction metric $G_{i}$, to balance quality and latency. The satisfaction metric $G_{i}$ of the task $i$ is denoted as
@@ IV Satisfaction: Quality Control for IIoT
In the context of combining the IIoT with FL, the final aggregated global model depends on the local models contributed by individual nodes, while the size of the raw data ${D_{i}}$ reflects the node’s contribution to the overall FL training process. ${D_{i}}$ is denoted as ${D_{i}}=\frac{T}{{{\theta_{i}}}}d$, where $d$ is the amount of data collected per unit time period.
@@ V-A Utility Model and Problem Formulation
To meet the quality and latency requirements of IIoT tasks, we formulate the corresponding utility functions and optimization objectives. Node $i$ is incentivized to handle incoming task requests, and each participating node receives a monetary reward ${R_{i}}$ from the server. Therefore, the utility of node $i$ is the difference between the reward ${R_{i}}$ and the cost ${C_{i}}$ of participating in the FL task. The utility $U_{i}$ can be expressed as
@@
In order to meet the quality and latency requirements of IIoT tasks, we define the goal of the incentive mechanism as maximizing server utility function $V$ and node utility function $U_{i}$. The optimization problem for the node utility is formulated as follows: